Find the probility that 2 persons will have


1.SAME BIRTHDAY


2.DIFFERENT BIRTHDAY.


(ignoring a leap year)

Grade:10

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

dear student,

eliminating the leap year case,

chance of a day being a person birthday is = 1/365

probability of both having same birthday is = 1/365^2

probability of having different birthdays = 1-(1/365^2)

all the best